# Hole burning experiments and modeling in erbium-doped silica glass fibers down to millikelvin temperatures: evidence for ultra-long population storage

### v2

@@ Abstract
We use spectral hole burning to investigate spin dynamics within the electronic Zeeman sublevels of the ground state, ${}^{4}I_{15/2}$, of the erbium (Er3+) ions in erbium-doped fibers (EDF). Conducted at ultra-low temperatures and under varying magnetic fields, our study reveals distinct changes in spin relaxation dynamics across different conditions. We identified three decay components at $\sim 7$ mK, with one achieving spin lifetimes of over 9 hours under optimal conditions, while two components were observed at higher temperatures. The fairly stable relative weights of the decay components across conditions suggest distinct ion populations contributing to the observed relaxation dynamics. While earlier studies struggled to account for all decay components at higher temperatures, our approach successfully models spin dynamics across all observed decay components, using a consistent set of underlying mechanisms—spin flip-flop interactions, direct coupling to two-level systems, and Raman-type processes—and distinguishing the decay components by the strengths with which these mechanisms contribute. These results suggest EDFs’ potential as a promising candidate for quantum memory applications, with further room for optimization.
@@ I Introduction
Recent advances in SHB studies of EDFs have revealed long-lived spin states within the Zeeman sublevels of Er3+ at low temperatures and under applied magnetic fields [14, 23, 4, 24]. In this study, we investigate the spectral hole properties of EDFs under a broad range of conditions, specifically focusing on the spin dynamics within Zeeman sublevels of the Er3+ ground state ${}^{4}I_{15/2}$. Our observations indicate narrower spectral holes than previously reported on REI-doped glasses [14, 3, 28, 29, 30], suggesting enhanced coherence properties and reduced environmental interactions at ultra-low temperatures. Notably, our experiments demonstrate spin lifetimes exceeding 9 hours at $\sim 7$ mK under optimal magnetic field conditions, suggesting the potential of EDFs for extended quantum state storage and promising applications in quantum communication networks. The stable weights of these decay components across various conditions suggest the existence of distinct ion classes, contributing uniquely to the relaxation dynamics.
@@ II Experimental Setup and Measurement
For optical excitation and scanning, we used a Toptica DL Pro continuous-wave (CW) laser operating at 1532 nm (see Fig. 1). The laser has a nominal linewidth below 1 MHz and shows excellent short-term stability during the burn process. Over longer acquisition timescales, the readout scans occasionally exhibited frequency drift up to $\sim 50$ MHz, which was corrected by centering each spectral hole during analysis. A 50/50 beamsplitter (BS) directed a portion of the laser beam to a wavemeter for real-time monitoring of wavelength and power. A polarizing beam splitter (PBS) controlled the polarization for effective spectral hole burning.
@@ III Results and Analysis
This section presents SHB experiments on EDF across a wide range of temperatures ($\sim$7 to 2400 mK) and magnetic fields (up to 200 mT), focusing on spin relaxation dynamics and spectral hole characteristics. Subsection III.1, which investigates spectral hole properties, is divided into two parts: Part III.1.1 explores temperature- and field-dependent changes in spectral hole behavior, including the emergence of additional decay components at ultra-low temperatures, while Part III.1.2 examines the stability of decay component weights across various conditions. Subsection III.2 describes how an established model represents spin relaxation rates in terms of temperature and magnetic field dependencies. Subsection III.3 applies the model to two- and three-component spin dynamics, using the same temperature and magnetic field dependencies across all components while allowing variations in the strengths of the mechanisms.
@@ III.1.1 Spectral Hole Behavior and Decay Components
As the temperature decreases below 80 mK, the spectral holes shift to Gaussian profiles, indicating a transition in the dominant broadening mechanism. At ultra-low temperatures, where phonon-driven decoherence is strongly suppressed, static inhomogeneities such as local strain or site disorder become more prominent, giving rise to Gaussian line shapes [33, 34]. At $\sim 7$ mK, we observed a spectral hole width of 7.5 MHz with a spin-transfer efficiency of 92.5% and a pronounced extension of erbium spin lifetimes by several orders of magnitude. This narrow spectral hole represents the narrowest linewidth reported for REI-doped glasses [14, 3, 28, 29, 30]. The Gaussian profile observed at $T=7$ mK, as shown in Figure 3b, highlights this transition and contrasts with the Lorentzian shapes characteristic of higher temperatures.
@@
At $\sim 7$ mK, however, a third decay component, $T_{c}$, is necessary to describe the dynamics accurately. The inadequacy of a two-component model at $\sim 7$ mK is evident in Figure 3a, in which the two-component fit (red dashed curve) fails to capture the full decay behavior. The inclusion of the third component, $T_{c}$, significantly improves the representation (green curve) of the decay, with $T_{c}$ exhibiting a characteristic time on the order of thousands of seconds (up to around 34,000 seconds), indicating much longer spin lifetimes. The inadequacy of a two-component model to fully capture this long-lived behavior is reflected in both the fit residuals and the persistence of the spectral hole over many hours. This ultra-slow decay is not a fitting artifact but represents a distinct and experimentally supported population dynamic that motivates the inclusion of a third exponential component.
@@ III.1.2 Decay Component Weights
Figure 4: Weights for two decay components at higher temperatures, with measurements across different conditions fitted to a constant to determine the average weight of each decay component. Panel (a) shows weights at varying temperatures for a fixed magnetic field ($B=50\,\text{mT}$), while panels (b) and (c) show weights across different magnetic fields at $T=800\,\text{mK}$ and $T=80\,\text{mK}$, respectively. The decay component weights $W_{a}=0.62$ and $W_{b}=0.38$ are shown with shaded regions representing the standard deviation (±$\pm 1\sigma$), highlighting the clustering of weights around their means, which suggests the existence of ion classes.
Figure 5: Weights for three decay components at $T=7\,\text{mK}$ across different magnetic fields. The weights were fitted to a constant, yielding average weights $W_{a}=0.52$, $W_{b}=0.26$, and $W_{c}=0.22$. The shaded regions represent the standard deviation (±$\pm 1\sigma$) for each decay component weight, illustrating the variability around their means.
@@
The weights of the two decay components exhibit a trend toward more equal contributions at higher temperatures, as seen in Figure 4a. Across varying magnetic fields, the weights also display relative stability, illustrated in Figures 4b and 4c. While some local fluctuations are observed, such as the near-crossover at 80 mT in Fig.4c and the gradual convergence of weights above $\sim$1.5 K in Fig.4a, these occur in regions where the extraction of decay weights is more susceptible to uncertainty. This includes the onset of the third decay component near 80 mK and increased noise or reduced contrast at higher temperatures. The error bars shown reflect statistical uncertainty from the fits and may not fully capture additional sources of variation, such as overlapping decay components or limited dynamic range. Although these fluctuations stand out relative to adjacent points, they remain limited in number, and the broader dataset supports the presence of distinct ion classes under most experimental conditions.
At $\sim 7$ mK, the emergence of a third component introduces a new distribution, with the decay component $A$ contributing an average weight $W_{a}=52\%$ (standard deviation: 7%), the decay component $B$ contributing an average weight $W_{b}=26\%$ (standard deviation: 5%), and the decay component $C$, associated with $T_{c}$ (longest-lived ions), contributing an average weight $W_{c}=22\%$ (standard deviation: 5%) to the total weight.
@@ III.2 Modeling of the Spin Lifetime Mechanism
The parameters $l$, $m$, and $n$ describe the field and temperature dependencies of the TLS and Raman terms and are the same across all decay components, providing a unified model for the underlying mechanisms. Each mechanism contributes additively to the total decay rate in Eq. 1, and the complete expressions for each term are constructed to yield results with units of $\mathrm{s}^{-1}$. The scaling parameters $\alpha^{i}_{\mathrm{ff}}$, $\alpha^{i}_{\mathrm{TLS}}$, and $\alpha^{i}_{\mathrm{R}}$ vary between decay components, capturing differences in the contributions of each mechanism for the different ion classes. At higher temperatures, we identify two distinct ion classes ($T_{a}$, $T_{b}$), while at low temperatures ($\sim 7\,\mathrm{mK}$), a third class ($T_{c}$) emerges. This framework allows the model to remain consistent across experimental conditions while accommodating the distinct relaxation behaviors observed in each ion class.
@@ III.3 Model Fitting to Experimental Data
In the low-temperature regime ($\sim 7$ mK), the model incorporates the third decay component ($T_{c}$), capturing the behavior of all three components ($T_{a}$, $T_{b}$, $T_{c}$) under varying magnetic fields, as shown in Figure 8. At this temperature, the Raman process ($\alpha^{i}_{\text{R}}=0$) becomes negligible, leaving direct coupling and flip-flop interactions as the only contributors. Notably, the contribution of direct coupling to TLS ($\alpha^{i}_{\text{TLS}}$) is significantly reduced for $T_{c}$ compared to $T_{a}$ and $T_{b}$, aligning with the much longer lifetime of $T_{c}$ at higher magnetic fields. This reduction in direct coupling to TLS for $T_{c}$ is a key factor contributing to its extended lifetime and may explain the emergence of this long-lived ion class at ultra-low temperatures.
@@
The extracted exponents ($l=1.35\pm 0.05$, $m=0.20\pm 0.02$, $n=3.0\pm 0.1$) provide a unified description of spin relaxation dynamics. The value of $l$ is slightly higher than that found in the previous study [14], while the value $n=3.0$ aligns with it. For $m$, the earlier study [14], which focused on the higher-temperature regime (down to 800 mK), reported a value of $m=1.2$, and when we fit only our higher-temperature data, we find that this value remains compatible with the observed behavior. However, to model the full dataset — including the ultra-low-temperature regime at $\sim 7$ mK and 80 mK — a reduced value of $m=0.20\pm 0.02$ was required. This lower exponent provides an accurate fit across all temperature regimes and maintains agreement with the high-temperature data as well. We therefore consider this updated value to be a more comprehensive and consistent description of the temperature dependence in our broader experimental range. The g-factor ($g=0.50\pm 0.05$) extracted in this study is notably small compared to some values reported for erbium ions [37, 38, 39, 40].
@@ IV Discussion and Outlook
This study presents an analysis of spin relaxation dynamics within the electronic Zeeman sublevels of the ground state ${}^{4}I_{13/2}$ of the erbium (Er3+) in erbium-doped fibers (EDF). The findings cover a wide temperature range ($\sim 7$ to 2400 mK) and magnetic field range (0 mT to 200 mT), revealing several notable insights that advance our understanding of these systems and their potential for quantum memory applications.
Our results reveal distinct spin relaxation dynamics across three temperature ranges: higher temperatures (above 80 mK), a transition range (44–80 mK), and ultra-low temperatures ($\sim 7$ mK).
@@
At $\sim 7$ mK, a third decay component $T_{c}$ emerged, with spin lifetimes exceeding 9 hours under optimal conditions. This exceptionally long lifetime aligns with the significantly reduced contribution of direct coupling to two-level systems (TLS) ($\alpha^{c}_{\text{TLS}}$) compared to $T_{a}$ and $T_{b}$. The suppression of this mechanism at ultra-low temperatures likely enables the distinct behavior of $T_{c}$, marking it as a long-lived ion class. The spectral holes at this temperature required Gaussian fits, indicating a transition in the broadening mechanism.
@@
The theoretical model used in this study, adapted from the prior work of [14], effectively captured the observed spin dynamics. While their study was limited to temperatures down to 650 mK and focused primarily on fastest decay components, our analysis extended to ultra-low temperatures, specifically down to 7 mK. The model accommodates three decay components at $\sim 7$ mK and two at other temperatures by employing shared temperature and magnetic field dependencies across all components while allowing the contributions of the underlying mechanisms to vary.
@@
Our results emphasize the potential of EDFs for quantum memory applications, particularly given the exceptionally long spin lifetimes exceeding 9 hours observed at $\sim 7$ mK. This suggests that EDFs may serve as stable, long-term quantum state storage systems, provided the ions corresponding to the longest decay component $T_{c}$ can be effectively isolated and utilized.

### v3

@@ Abstract
We use spectral hole burning to investigate spin dynamics within the electronic Zeeman sublevels of the ground state, $^{4}I_{15/2}$, of the erbium (Er3+) ions in erbium-doped fibers (EDF). Conducted at ultra-low temperatures and under varying magnetic fields, our study reveals distinct changes in spin relaxation dynamics across different conditions. We identified three decay components at approximately 7 mK, with one achieving spin lifetimes of over 9 h under optimal conditions, while two components were observed at higher temperatures. The fairly stable relative weights of the decay components across conditions suggest distinct ion populations contributing to the observed relaxation dynamics. While earlier studies struggled to account for all decay components at higher temperatures, our approach successfully models spin dynamics across all observed decay components, using a consistent set of underlying mechanisms—spin flip-flop interactions, direct coupling to two-level systems, and Raman-type processes—and distinguishing the decay components by the strengths with which these mechanisms contribute. These results suggest EDFs’ potential as a promising candidate for quantum memory applications, with further room for optimization.
@@ I Introduction
Recent advances in SHB studies of EDFs have revealed long-lived spin states within the Zeeman sublevels of Er3+ at low temperatures and under applied magnetic fields [14, 23, 4, 24]. In this study, we investigate the spectral hole properties of EDFs under a broad range of conditions, specifically focusing on the spin dynamics within Zeeman sublevels of the Er3+ ground state $^{4}I_{15/2}$. Our observations indicate narrower spectral holes than previously reported on REI-doped glasses [14, 3, 28, 5, 29], suggesting enhanced coherence properties and reduced environmental interactions at ultra-low temperatures. Notably, our experiments demonstrate spin lifetimes exceeding 9 hours at approximately 7 mK under optimal magnetic field conditions, suggesting the potential of EDFs for extended quantum state storage and promising applications in quantum communication networks. The stable weights of these decay components across various conditions suggest the existence of distinct ion classes, contributing uniquely to the relaxation dynamics.
@@ II Experimental Setup and Measurement
For optical excitation and scanning, we used a Toptica DL Pro continuous-wave (CW) laser operating at 1532 nm (see Fig. 1). The laser has a nominal linewidth below 1 MHz and shows excellent short-term stability during the burn process. Over longer acquisition timescales, the readout scans occasionally exhibited frequency drift up to approximately 50 MHz, which was corrected by centering each spectral hole during analysis. A 50/50 beamsplitter (BS) directed a portion of the laser beam to a wavemeter for real-time monitoring of wavelength and power. A polarizing beam splitter (PBS) controlled the polarization for effective spectral hole burning.
@@ III Results and Analysis
This section presents SHB experiments on EDF across a wide range of temperatures (approximately 7 to 2400 mK) and magnetic fields (up to 200 mT), focusing on spin relaxation dynamics and spectral hole characteristics. Subsection III.1, which investigates spectral hole properties, is divided into two parts: Part III.1.1 explores temperature- and field-dependent changes in spectral hole behavior, including the emergence of additional decay components at ultra-low temperatures, while Part III.1.2 examines the stability of decay component weights across various conditions. Subsection III.2 describes how an established model represents spin relaxation rates in terms of temperature and magnetic field dependencies. Subsection III.3 applies the model to two- and three-component spin dynamics, using the same temperature and magnetic field dependencies across all components while allowing variations in the strengths of the mechanisms.
@@ III.1.1 Spectral Hole Behavior and Decay Components
As the temperature decreases below 80 mK, the spectral holes shift to Gaussian profiles, indicating a transition in the dominant broadening mechanism. At ultra-low temperatures, where phonon-driven decoherence is strongly suppressed, static inhomogeneities such as local strain or site disorder become more prominent, giving rise to Gaussian line shapes [32, 33]. At approximately 7 mK, we observed a spectral hole width of 7.5 MHz with a spin-transfer efficiency of 92.5% and a pronounced extension of erbium spin lifetimes by several orders of magnitude. This narrow spectral hole represents the narrowest linewidth reported for REI-doped glasses [14, 3, 28, 5, 29]. The Gaussian profile observed at $T=7$ mK, as shown in Figure 3b, highlights this transition and contrasts with the Lorentzian shapes characteristic of higher temperatures.
@@
At approximately 7 mK, however, a third decay component, $T_{c}$, is necessary to describe the dynamics accurately. The inadequacy of a two-component model at approximately 7 mK is evident in Figure 3a, in which the two-component fit (red dashed curve) fails to capture the full decay behavior. The inclusion of the third component, $T_{c}$, significantly improves the representation (green curve) of the decay, with $T_{c}$ exhibiting a characteristic time on the order of thousands of seconds (up to around 34,000 seconds), indicating much longer spin lifetimes. The inadequacy of a two-component model to fully capture this long-lived behavior is reflected in both the fit residuals and the persistence of the spectral hole over many hours. This ultra-slow decay is not a fitting artifact but represents a distinct and experimentally supported population dynamic that motivates the inclusion of a third exponential component.
@@ III.1.2 Decay Component Weights
Figure 4: Weights for two decay components at higher temperatures, with measurements across different conditions fitted to a constant to determine the average weight of each decay component. Panel (a) shows weights at varying temperatures for a fixed magnetic field ($B=50\,\text{mT}$), while panels (b) and (c) show weights across different magnetic fields at $T=800\,\text{mK}$ and $T=80\,\text{mK}$, respectively. The decay component weights $W_{a}=0.62$ and $W_{b}=0.38$ are shown with shaded regions representing the standard deviation ($\pm 1\sigma$), highlighting the clustering of weights around their means, which suggests the existence of ion classes.
Figure 5: Weights for three decay components at $T=7\,\text{mK}$ across different magnetic fields. The weights were fitted to a constant, yielding average weights $W_{a}=0.52$, $W_{b}=0.26$, and $W_{c}=0.22$. The shaded regions represent the standard deviation ($\pm 1\sigma$) for each decay component weight, illustrating the variability around their means.
@@
The weights of the two decay components exhibit a trend toward more equal contributions at higher temperatures, as seen in Figure 4a. Across varying magnetic fields, the weights also display relative stability, illustrated in Figures 4b and 4c. While some local fluctuations are observed, such as the near-crossover at 80 mT in Fig.4c and the gradual convergence of weights above approximately 1.5 K in Fig.4a, these occur in regions where the extraction of decay weights is more susceptible to uncertainty. This includes the onset of the third decay component near 80 mK and increased noise or reduced contrast at higher temperatures. The error bars shown reflect statistical uncertainty from the fits and may not fully capture additional sources of variation, such as overlapping decay components or limited dynamic range. Although these fluctuations stand out relative to adjacent points, they remain limited in number, and the broader dataset supports the presence of distinct ion classes under most experimental conditions.
At approximately 7 mK, the emergence of a third component introduces a new distribution, with the decay component $A$ contributing an average weight $W_{a}=52\%$ (standard deviation: 7%), the decay component $B$ contributing an average weight $W_{b}=26\%$ (standard deviation: 5%), and the decay component $C$, associated with $T_{c}$ (longest-lived ions), contributing an average weight $W_{c}=22\%$ (standard deviation: 5%) to the total weight.
@@ III.2 Modeling of the Spin Lifetime Mechanism
The parameters $l$, $m$, and $n$ describe the field and temperature dependencies of the TLS and Raman terms and are the same across all decay components, providing a unified model for the underlying mechanisms. Each mechanism contributes additively to the total decay rate in Eq. 1, and the complete expressions for each term are constructed to yield results with units of $\mathrm{s}^{-1}$. The scaling parameters $\alpha^{i}_{\mathrm{ff}}$, $\alpha^{i}_{\mathrm{TLS}}$, and $\alpha^{i}_{\mathrm{R}}$ vary between decay components, capturing differences in the contributions of each mechanism for the different ion classes. At higher temperatures, we identify two distinct ion classes ($T_{a}$, $T_{b}$), while at low temperatures (approximately 7 mK), a third class ($T_{c}$) emerges. This framework allows the model to remain consistent across experimental conditions while accommodating the distinct relaxation behaviors observed in each ion class.
@@ III.3 Model Fitting to Experimental Data
In the low-temperature regime (approximately 7 mK), the model incorporates the third decay component ($T_{c}$), capturing the behavior of all three components ($T_{a}$, $T_{b}$, $T_{c}$) under varying magnetic fields, as shown in Figure 8. At this temperature, the Raman process ($\alpha^{i}_{\text{R}}=0$) becomes negligible, leaving direct coupling and flip-flop interactions as the only contributors. Notably, the contribution of direct coupling to TLS ($\alpha^{i}_{\text{TLS}}$) is significantly reduced for $T_{c}$ compared to $T_{a}$ and $T_{b}$, aligning with the much longer lifetime of $T_{c}$ at higher magnetic fields. This reduction in direct coupling to TLS for $T_{c}$ is a key factor contributing to its extended lifetime and may explain the emergence of this long-lived ion class at ultra-low temperatures.
@@
The extracted exponents ($l=1.35\pm 0.05$, $m=0.20\pm 0.02$, $n=3.0\pm 0.1$) provide a unified description of spin relaxation dynamics. The value of $l$ is slightly higher than that found in the previous study [14], while the value $n=3.0$ aligns with it. For $m$, the earlier study [14], which focused on the higher-temperature regime (down to 800 mK), reported a value of $m=1.2$, and when we fit only our higher-temperature data, we find that this value remains compatible with the observed behavior. However, to model the full dataset — including the ultra-low-temperature regime at approximately 7 mK and 80 mK — a reduced value of $m=0.20\pm 0.02$ was required. This lower exponent provides an accurate fit across all temperature regimes and maintains agreement with the high-temperature data as well. We therefore consider this updated value to be a more comprehensive and consistent description of the temperature dependence in our broader experimental range. The g-factor ($g=0.50\pm 0.05$) extracted in this study is notably small compared to some values reported for erbium ions [36, 37, 38, 39].
@@ IV Discussion and Outlook
This study presents an analysis of spin relaxation dynamics within the electronic Zeeman sublevels of the ground state $^{4}I_{13/2}$ of the erbium (Er3+) in erbium-doped fibers (EDF). The findings cover a wide temperature range (approximately 7 to 2400 mK) and magnetic field range (0 mT to 200 mT), revealing several notable insights that advance our understanding of these systems and their potential for quantum memory applications.
Our results reveal distinct spin relaxation dynamics across three temperature ranges: higher temperatures (above 80 mK), a transition range (44–80 mK), and ultra-low temperatures (approximately 7 mK).
@@
At approximately 7 mK, a third decay component $T_{c}$ emerged, with spin lifetimes exceeding 9 hours under optimal conditions. This exceptionally long lifetime aligns with the significantly reduced contribution of direct coupling to two-level systems (TLS) ($\alpha^{c}_{\text{TLS}}$) compared to $T_{a}$ and $T_{b}$. The suppression of this mechanism at ultra-low temperatures likely enables the distinct behavior of $T_{c}$, marking it as a long-lived ion class. The spectral holes at this temperature required Gaussian fits, indicating a transition in the broadening mechanism.
@@
The theoretical model used in this study, adapted from the prior work of [14], effectively captured the observed spin dynamics. While their study was limited to temperatures down to 650 mK and focused primarily on fastest decay components, our analysis extended to ultra-low temperatures, specifically down to 7 mK. The model accommodates three decay components at approximately 7 mK and two at other temperatures by employing shared temperature and magnetic field dependencies across all components while allowing the contributions of the underlying mechanisms to vary.
@@
Our results emphasize the potential of EDFs for quantum memory applications, particularly given the exceptionally long spin lifetimes exceeding 9 hours observed at approximately 7 mK. This suggests that EDFs may serve as stable, long-term quantum state storage systems, provided the ions corresponding to the longest decay component $T_{c}$ can be effectively isolated and utilized.